# Why Does New Knowledge Create Messy Ripple Effects in LLMs?

### v2

@@ Abstract
Extensive previous research has focused on post-training knowledge editing (KE) for language models (LMs) to ensure that knowledge remains accurate and up-to-date. One desired property and open question in KE is to let edited LMs correctly handle ripple effects, where LM is expected to answer its logically related knowledge accurately. In this paper, we answer the question of why most KE methods still create messy ripple effects. We conduct extensive analysis and identify a salient indicator, GradSim, that effectively reveals when and why updated knowledge ripples in LMs. GradSim is computed by the cosine similarity between gradients of the original fact and its related knowledge. We observe a strong positive correlation between ripple effect performance and GradSim across different LMs, KE methods, and evaluation metrics. Further investigations into three counter-intuitive failure cases (Negation, Over-Ripple, Multi-Lingual) of ripple effects demonstrate that these failures are often associated with very low GradSim. This finding validates that GradSim is an effective indicator of when knowledge ripples in LMs.
@@
Jiaxin Qin1,2, Zixuan Zhang [^1] footnotemark: 1, Chi Han1, Manling Li1, Pengfei Yu1, Heng Ji1 1University of Illinois Urbana-Champaign 2Renmin University of China {qjx0814, zixuan11, chihan3, pengfei4, hengji}@illinois.edu manling@stanford.edu
@@ 1 Introduction and Related Work
Large language models (LLMs) can serve as powerful knowledge bases (KBs) thanks to their impressive knowledge storage, retrieval, and reasoning capabilities (Petroni et al., 2019; AlKhamissi et al., 2022). However, real-world knowledge keeps updating and evolving constantly, which motivates extensive research efforts on post-training knowledge editing (KE) Meng et al. (2022, 2023); Yin et al. (2023); Zhong et al. (2023); Song et al. (2024); Liu et al. (2024) to make sure that the knowledge in LMs remains accurate and up-to-date.
@@ 2 GradSim: A Ripple Effect Indicator
$\textit{GradSim}(x,y)=\cos(\nabla_{\theta}P_{\theta}(a_{x}|q_{x}),\nabla_{\theta}P_{\theta}(a_{y}|q_{y})).$
@@ 4 Conclusion
Through extensive experiments and analysis, we propose GradSim, computed as the cosine similarity between gradients of the original fact and its related knowledge, as a crucial indicator for the effectiveness of the ripple effects. The positive correlation observed between GradSim and ripple effect performance across various LMs, KE methods, and evaluation metrics underscores its reliability. Additionally, our exploration of failure cases such as Negation, Over-Ripple, and Cross-Lingual scenarios, further confirms that low GradSim values are indicative of ripple effect failures.
@@
Appendix A Extra Results of GradSim
The RippleEdits dataset comprises six distinct tasks: Logical Generalization (LG), Compositionality I (CI), Compositionality II (CII), Subject Aliasing (SA), Preservation (PV), and Relation Specificity (RS) (Cohen et al., 2023). These tasks are designed to evaluate different aspects of knowledge editing in neural networks. Specifically, LG, CI, CII, and SA are tasks where the model is expected to manifest new knowledge in response to edits made to existing entries. Conversely, in the RS and PV tasks, the existing knowledge should remain unaltered post-editing, as these tasks are designed to test the model’s ability to preserve information that is logically independent from the changes applied.
@@ Appendix A Extra Results of GradSim
Appendix B How to Represent Knowledge Distribution?
@@
B.1 Does the way that we express a piece of knowledge change the knowledge distribution?

### v3

@@ Abstract
Extensive previous research has focused on post-training knowledge editing (KE) for language models (LMs) to ensure that knowledge remains accurate and up-to-date. One desired property and open question in KE is to let edited LMs correctly handle ripple effects, where LM is expected to answer its logically related knowledge accurately. In this paper, we answer the question of why most KE methods still create messy ripple effects. We conduct extensive analysis and identify a salient indicator, GradSim, that effectively reveals when and why updated knowledge ripples in LMs. GradSim is computed by the cosine similarity between gradients of the original fact and its related knowledge. We observe a strong positive correlation between ripple effect performance and GradSim across different LMs, KE methods, and evaluation metrics. Further investigations into three counter-intuitive failure cases (Negation, Over-Ripple, Multi-Lingual) of ripple effects demonstrate that these failures are often associated with very low GradSim. This finding validates that GradSim is an effective indicator of when knowledge ripples in LMs. The code is available.22footnotemark: 2
@@
Jiaxin Qin1111Work done during internship at UIUC., Zixuan Zhang1, Pengfei Yu1,3, Manling Li2, Heng Ji1 1University of Illinois Urbana-Champaign 2Stanford University 3Boson AI {qjx0814, zixuan11, pengfei4, hengji}@illinois.edu manlingl@stanford.edu
@@ 1 Introduction and Related Work
Large language models (LLMs) can serve as powerful knowledge bases (KBs) thanks to their impressive knowledge storage, retrieval, and reasoning capabilities (Petroni et al., 2019; AlKhamissi et al., 2022; Zhang et al., 2024). However, real-world knowledge keeps updating and evolving constantly, which motivates extensive research efforts on post-training knowledge editing (KE) Meng et al. (2022, 2023); Yin et al. (2023); Zhong et al. (2023a); Song et al. (2024); Liu et al. (2024) to make sure that the knowledge in LMs remains accurate and up-to-date.
@@ 2 GradSim: A Ripple Effect Indicator
$\textit{GradSim}(x,y)$ $=$
$\cos(\nabla_{\theta}P_{\theta}(a_{x}|q_{x}),$ $\nabla_{\theta}P_{\theta}(a_{y}|q_{y}))$
@@ 4 Conclusion
Through extensive experiments and analysis, we propose GradSim, computed as the cosine similarity between gradients of the original fact and its related knowledge, as a crucial indicator for the effectiveness of the ripple effects. The positive correlation observed between GradSim and ripple effect performance across various LMs, KE methods, and evaluation metrics underscores its reliability. Additionally, our exploration of failure cases further confirms that low GradSim values are indicative of ripple effect failures.
@@
Acknowledgement
This research is based upon work supported by U.S. DARPA SemaFor Program No. HR001120C0123, DARPA Seedling BRIES HR0011-24-3-0325, DARPA INCAS Program No. HR001121C0165, and DARPA ITM Program No. FA8650-23-C-7316. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
Appendix A Knoweledge Editing Methods
Current knowledge editing methods can be categorized into several distinct approaches, including fine-tuning-based methods, locate-then-edit methods, and meta-learning methods. In this paper, we employ a range of knowledge editing techniques to validate our theorem. A brief introduction to each of these methods will be provided here.
ROME: Rank-One Model Editing (ROME) conceptualizes an MLP module as a key-value store. In this framework, the key represents an encoded subject, while the value represents the knowledge associated with that subject. The MLP retrieves the corresponding value by accessing the key. ROME modifies the MLP weights using a rank-one adjustment to directly insert new key-value pairs.
MEMIT: Mass Editing Memory in a Transformer (MEMIT) builds upon ROME, and is designed to handle large-scale edits by inserting multiple memory entries simultaneously through modifications to the MLP weights across several key layers. MEMIT employs causal tracing to identify a set of mediating MLP layers that store and recall memories related to a specific subject. For a set of new memories, an update ($\Delta$) is computed and propagated across all the identified mediating MLP layers, ensuring that by the final layer, the output captures and reflects all the newly inserted memories.
MEND: Model Editor Networks with Gradient Decomposition (MEND) consist of small auxiliary networks designed to make quick, localized edits to a pre-trained model’s behavior using a single input-output pair. MEND achieves this by learning how to transform the gradient generated through standard fine-tuning, employing a low-rank decomposition of the gradient to make the transformation more computationally feasible. MEND can be trained efficiently on a single GPU in less than a day, even for models with over 10 billion parameters. Once trained, MEND allows for rapid application of new edits to the pre-trained model.
Appendix B Additional Experiments
B.1 Experiments with various knowledge experiments
To support our theorem, we also did experiments with more knowledge editing methods, including Fine-Tuning(FT) and MEND. In the basic FT process, we use Adam and early stopping to minimize the loss of new edits on full parameters. In the additional experiments, we only compute on GPT-XL due to computational resource limitations. The experiments are done with the popular subset in RippleEdits. Here are the experiment results:
Method Spearman Correlation Pearson Correlation Best Fit Line
MEND 0.628 0.501 $y=0.845x-0.177$
FT 0.231 0.731 $y=0.430x+0.119$
Table 1: Main experiments with other knowledge editing method
These experimental results demonstrate the versatility of the GradSim indicator.
B.2 Experiments on other dataset
We conducted experiments on all data in RippleEdits to demonstrate the indicator’s effectiveness. Additional experiments are also done on the 2-hop subset of MQUAKE dataset Zhong et al. (2023b) to support the validity of our result. The results are as follows
Method Model Spearman Correlation Pearson Correlation Best Fit Line
ROME GPT2-XL 0.767 0.783 $y=1.164x-8.923$
Table 2: Results on MQUAKE
Appendix C Extra Results of GradSim
The RippleEdits dataset comprises six distinct tasks: Logical Generalization (LG), Compositionality I (CI), Compositionality II (CII), Subject Aliasing (SA), Preservation (PV), and Relation Specificity (RS) (Cohen et al., 2023). These tasks are designed to evaluate different aspects of knowledge editing in neural networks. Specifically, LG, CI, CII, and SA are tasks where the model is expected to manifest new knowledge in response to edits made to existing entries. Conversely, in the RS and PV tasks, the existing knowledge should remain unaltered post-editing, as these tasks are designed to test the model’s ability to preserve information that is logically independent of the changes applied.
@@ Appendix C Extra Results of GradSim
Appendix D How to Represent Knowledge Distribution?
@@
D.1 Does the way that we express a piece of knowledge change the knowledge distribution?